OpenStudy (anonymous):

A rectangular prism has a surface area of 268 square centimeters. If the length is 8 cm and the width is 10 cm, what is the height of the prism?

OpenStudy (kropot72):

Area of top of prism = 8 * 10 square centimeters Area of bottom of prism = 8 * 10 square centimeters Let the height of the prism = h centimeters Area of sides of prism = (2 * 8 * h) + (2 * 10 * h) square centimeters Total surface area of prism = 268 = (2 * 8 * 10) + (2 * 8 * h) + (2 * 10 * h) square centimeters 268 = 160 + 16h + 20h ...............(1) Can you solve equation (1) to find the value of h ?